# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v2

@@ III-A Motivation and Rationale
III-B Target Synthesis and Composition
The DenseSIRST dataset, a blend of real and synthetic infrared images, provides a challenging and dense environment for training robust detection algorithms. We curated this dataset from diverse real-world scenarios, supplemented with synthetic targets designed for physical plausibility and visual realism.
The dataset leverages our SIRST V2 dataset, creating a broad library of infrared small targets. Each image is meticulously designed to emulate the densely populated surveillance scenarios. Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image. These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library. The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
To ensure a smooth integration of targets and their backgrounds, we employ a Gaussian-weighted copy-paste technique, an evolution of the concept from mixup. This modification is governed by the following equation:
$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
where $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location. Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target. $B_{box}$ represents the original target appearance. The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
The matrix $G$ denotes a Gaussian blur, employed during target pasting, and is defined as:
where the relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$. These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size. The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$. This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance. To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
where $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
III-C Dataset Description and Characteristics
III-C1 Dataset Composition
@@ III-C2 Annotations and Labels
III-C3 Statistical Characteristics
The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers an analysis of the dataset’s key characteristics.
Target Size Distribution: Most targets, as indicated in Fig. 3 (b), are extremely small, predominantly below $5\times 5$ pixels. The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery.
Local Contrast Distribution: The 90% targets have local contrast below $2$, as shown in Fig. 3 (c). This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
Target Brightness Distribution: Fig. 3 (d) reveals that only a minority of targets are the brightest points in their respective images. This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection.
In conclusion, the DenseSIRST dataset, with its wide range of image dimensions, prevalence of small target sizes, and low local contrast, provides a realistic and challenging benchmark for infrared small target detection algorithms.
@@ IV-A Overall Architecture and Training Pipeline
Figure 4: Schematic diagram of the proposed BAFE-Net network. BAFE-Net contains three modules: ResNet, FPN and BAFE-Head.
Figure 5: First, a dense area is randomly selected from the sky area of the image, and then several small targets are randomly selected from the small target library. After determining the appropriate pasting position, the Gaussian technique is used to integrate the small targets into the new background.
The overall architecture of BAFE-Net, which adopts the fully convolutional one-stage object detector (FCOS) [37] as its host network, is illustrated in Fig. 4. In addition to the original detection head present in the baseline, BAFE-Net introduces a segmentation head within the BAFE-Head module. The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes. By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information. Furthermore, BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to facilitate the bidirectional flow of information between the detection and segmentation branches. This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
To address the challenge of limited target samples, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP). This technique strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, ensuring that the pasted targets are only placed in sky regions. The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm. By augmenting the training dataset with realistically integrated synthetic targets, BAG-CP allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds. This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
@@ IV-B Dynamic Cross-Task Feature Exchange Head
For infrared small target detection, the scarcity of intrinsic target features and the abundance of similar background distractors pose significant challenges. To address this, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. The motivation behind this cross-task feature exchange is twofold:
  1. 1.
    Complementary Nature of Tasks: Although target detection and background segmentation have different objectives, they share a common goal of understanding the scene. By exchanging features between these tasks, each task can benefit from the complementary information learned by the other, leading to a more comprehensive understanding of the image.
  2. 2.
    Synergistic Relationship: Target detection focuses on identifying the presence and location of small targets, while background segmentation emphasizes the contextual information and local details of the scene. By sharing features between these tasks, target detection can leverage the semantic understanding of the background to better distinguish targets from distractors, while background segmentation can utilize the localized target information to refine its segmentation boundaries.
IV-B1 Dynamic Channel Selection
To further optimize cross-task feature exchange, we propose a dynamic selection mechanism that determines the most relevant feature channels for exchange based on their importance to task performance. The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution.
@@ IV-B1 Dynamic Channel Selection
Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we first compute a channel-wise attention weight vector $\mathbf{m}\in\mathbb{R}^{C}$ by a mechanism inspired by the design of our previous work Large Selective Kernel Network (LSKNet) [38]. This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion. The process can be summarized as follows:
$\mathbf{m}$ $=\sigma\left(\mathbf{X}\odot\left(\mathcal{F}(\mathbf{X})\odot\sigma\left(\mathcal{S}\left(\text{avg}(\mathbf{X})\oplus\max(\mathbf{X})\right)\right)\right)\right),$ (4)
@@
where $\odot$ denotes element-wise multiplication, $\oplus$ denotes the concatenation operation, and $\sigma$ is the Sigmoid function. $\max$ and avg are the maximum and average pooled spatial feature descriptors. $\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features. $\mathcal{F}$ is also a convolution operation applied to $\mathbf{X}$ , extracting local features.$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels. Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
@@
where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample. ${I}_{\text{topk}}\in\mathbb{R}^{K}$ represents the corresponding channel indices for these weights. This process dynamically selects the channels for exchange, improving the effectiveness of cross-task feature sharing.
IV-B2 Cross-Task Feature Hard-Exchange
Within BAFE-Net, we introduce a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Feature hard-exchange allows for the direct transfer of learned features from one task to another. By swapping entire feature channels, the receiving task has access to the explicit representation learned by the other task, facilitating a more direct integration of complementary information.
To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width. We define an exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
where $i$ represents the channel index. The exchange mask $\mathbf{e}$ is then expanded to the shape of $(C,H,W)$:
$\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (7)
where $\otimes$ denotes the expansion operation, and $\mathbf{1}_{H\times W}$ is a tensor of ones with a shape of $(H,W)$. $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ can obtain $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively through Eq. (4) to Eq. (7). The output tensors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ after the hard feature exchange are computed as:
IV-C Background-Aware Gaussian Copy-Paste
Infrared small target detection is inherently challenged by the paucity of intrinsic visual features in targets and the interference from similar background elements. Traditional data augmentation strategies, such as random cropping or flipping, are not effective in addressing these unique challenges. Consequently, we propose a novel data augmentation technique, Background-Aware Gaussian Copy-Paste (BAG-CP), which capitalizes on contextual information to enhance the training process. BAG-CP selectively augments training images by pasting infrared small targets into semantically relevant background areas, generating more realistic and diverse training samples. The fundamental idea is to bolster the detector’s ability to discern real targets from background noise through the seamless blending of pasted targets into their new environment. As illustrated in Fig. 5, The BAG-CP process unfolds as follows:
Prerequisite: Semantic Annotation of Background: Before performing the copy-paste operation, it is crucial to have the image segmented into different semantic regions. The DenseSIRST dataset includes annotated background categories, providing detailed background masks that separate sky from non-sky regions.
Step 1: Selection and Extraction of Infrared Small Targets: A library of small targets is created during the generation of the DenseSIRST dataset. A random selection of small targets from this library is made for pasting into the current training image.
Step 2: Gaussian Smoothing of Edges: Prior to pasting, the selected targets are subjected to Gaussian smoothing. This step ensures a smooth transition around the pasted targets, minimizing any conspicuous artifacts that might disrupt the learning process.
Step 3: Context-Aware Target Pasting Based on Background Semantics: The core of BAG-CP lies in context-aware placement. Using the semantic annotation of the background, we classify it into sky and non-sky regions. When choosing where to paste the Gaussian-smoothed targets, we only select locations within the sky regions. This approach mimics real-world scenarios where infrared small targets, especially aerial targets, are more likely to appear against a sky background.
During the pasting process, 1-3 regions are selected within the image to form clusters of small targets. Each region is 20 × 20 pixels in size and is populated with 8 to 12 targets randomly chosen from the target library. The selected targets are resized to within 5 × 5 pixels to maintain proportional variability and are pasted with a spacing of 1 to 2 pixels to simulate realistic clustering. The Gaussian-weighted edge blurring technique used follows Eq. (1), Eq. (2), and Eq. (3).
V Experiments
In this section, we outline a series of experimental evaluations designed to validate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method and the Background-Aware Feature Exchange Network (BAFE-Net) proposed in this study. Additionally, we conduct comparative analyses with state-of-the-art methods in infrared small target detection, using both quantitative metrics and qualitative assessments. We address the following pivotal questions in our experimental analysis:
  1. 1.
    Ablation Study on BAG-CP. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
  2. 2.
    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlight the contributions of each component in BAFE-Net.
  3. 3.
    Feature Exchange Mechanism Analysis. We investigate the impact of different feature exchange strategies on the performance of BAFE-Net to determine the optimal approach for exchanging features between the target detection head and the background segmentation head.
  4. 4.
    Impact of Backbone Depth: We examine the influence of varying backbone depths on the performance of the BAFE-Net. The primary aim is to establish the optimal depth that maximizes target detection while balancing computational efficiency.
  5. 5.
    Comparative Analysis with State-of-the-Art Methods. We conduct a comprehensive comparison between our method and other state-of-the-art methods on the DenseSIRST dataset.
V-A Experimental Settings
1) Dataset: During the experimental phase of our research, we conducted tests on the proposed dataset DenseSIRST.
2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch. We choose DAdaptAdam as the optimizer, learning rate is set to 1.0, and weight decay is set to 0.05. Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework.
@@ V-A Experimental Settings
TABLE II: Comparison with Other State-of-the-art methods on DenseSIRST.
Method Backbone DenseSIRST
FLOPs $\downarrow$ Params $\downarrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
              One-stage
SSD [39] 87.552G 23.746M 0.211 0.421 0.178 0.424
RetinaNet [40] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
YOLOv3 [41] 50.002G 61.949M 0.233 0.424 0.207 0.413
CenterNet [42] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
FCOS [37] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
ATSS [43] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
CentripetalNet [44] 0.491T 0.206G 0.244 0.259 0.201 0.244
AutoAssign [45] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
GFL [46] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
PAA [47] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
VFNet [48] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
PVT-T [49] 41.623G 21.325M 0.109 0.481 0.093 0.501
YOLOF [50] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
YOLOX [51] 8.578G 8.968M 0.210 0.341 0.180 0.331
TOOD [12] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
DyHead [52] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
DDOD [53] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
RTMDet [54] 51.278G 52.316M 0.229 0.349 0.212 0.350
EfficientDet [55] 34.686G 18.320M 0.146 0.464 0.094 0.517

### v3

@@ III-A4 Geometric and Appearance Augmentation
III-B Dataset Description and Characteristics
III-B1 Dataset Composition
@@ III-B2 Annotations and Labels
III-B3 Density Definition
To quantitatively capture the degree of target clustering, we define a density measure tailored for cluster regions within the dataset. A cluster region is identified when the Euclidean distance between the centroids of any two targets is less than 10 pixels, representing a meaningful spatial proximity threshold for dense grouping.
Formally, the density $\mathcal{D}$ of a cluster region is computed as:
$\mathcal{D}=\mathcal{N}\times\frac{\mathcal{A}_{t}}{\mathcal{A}_{c}},$ (4)
where $\mathcal{N}$ denotes the number of targets within the cluster, $\mathcal{A}{t}$ is the cumulative pixel area covered by all target masks inside the cluster, and $\mathcal{A}{c}$ represents the total pixel area of the cluster region itself. This metric jointly reflects both the quantity of targets and their relative spatial occupancy, providing a nuanced descriptor of local target density.
For consistent comparison across images, density values are normalized via min-max scaling based on the global minimum and maximum densities observed in both clustered and isolated target regions. Empirical statistics reveal a pronounced density difference, cluster regions exhibit an average density of 0.4092 compared to 0.1342 in single target regions, underscoring the metrics effectiveness in distinguishing densely grouped targets from sparse ones.
III-B4 Statistical Characteristics
Fig. 3 illustrates key statistical properties of the DenseSIRST dataset, emphasizing the detection challenges it entails.
  • Target Size Distribution (Fig. 3 (a)): The vast majority of targets are extremely small, predominantly under $5\times 5$ pixels, with the most frequent size being $3\times 3$ pixels. This distribution aligns with the typical scale of small infrared targets.
  • Local Contrast Distribution (Fig. 3 (b)): Approximately 90% of targets possess a local contrast below 2 relative to their surrounding background. This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
  • Target Brightness Distribution (Fig. 3 (c)): The brightness values of targets span a wide spectrum, with only a minority being the brightest points in their respective images. This diversity invalidates brightness as a reliable standalone feature, highlighting the necessity for robust algorithms that leverage richer contextual and structural information.
In conclusion, the DenseSIRST dataset exhibits rich annotation modalities, diverse environmental scenarios, and statistically challenging target characteristics. These attributes collectively establish it as a comprehensive and realistic benchmark, particularly well suited for developing and evaluating infrared small target detection algorithms under dense and low-contrast conditions.
@@ IV-A Overall Architecture and Training Pipeline
Figure 4: The framework of BAFE-Net. BAFE-Net contains three modules: ResNet, FPN and BAFE-Head. The Head module integrates a background segmentation branch to work in parallel with the classification branch. Both branches utilize the DCS Module to select the most discriminative top-k channel features, followed by channel-wise feature interaction. This enhancement transforms the original single-task object detection framework into a multi-task learning architecture, enabling simultaneous and efficient execution of both object detection and background semantic segmentation.
The proposed BAFE-Net architecture, which builds upon the fully convolutional one-stage object detector (FCOS) [22], is illustrated in Fig. 4. The network comprises three key components: ResNet backbone, Feature Pyramid Network (FPN), and BAFE-Head module. The BAFE-Head features two parallel branches: the original FCOS detection branch and an innovative background segmentation branch. This dual-branch design enables explicit modeling of background semantics, which is essential for distinguishing real targets from background clutter in infrared scenes.
BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism that facilitates bidirectional information flow between the detection and segmentation branches. Both branches incorporate a Dynamic Channel Selection (DCS) module, which identifies and utilizes the top-k most discriminative channel features to enable efficient Channel-wise Feature Exchange. This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
To address the challenge of limited target samples, BAFE-Net adopts the dataset synthesis method BAG-CP as data augmentation strategy. By leveraging semantic context, this method pastes synthetic infrared small targets exclusively onto sky regions in real-scene backgrounds, with Gaussian smoothing applied to the edges for seamless integration. This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
@@ IV-B Dynamic Cross-Task Feature Exchange Head
Infrared small target detection presents inherent challenges due to the weak discriminative characteristics of targets and the prevalence of complex background clutter. Conventional single-branch detection frameworks often struggle to effectively distinguish true targets from background distractors, particularly in scenarios characterized by limited target information and ambiguous contextual cues.
To address these challenges, we propose a multi-task collaborative framework that integrates object detection and background semantic segmentation within a unified prediction head, referred to as BAFE-Head. The core motivation behind this design lies in the complementary nature of the two tasks: object detection is tailored to localize small, sparse targets, whereas semantic segmentation captures global scene semantics and background structures. By jointly optimizing these tasks, the proposed framework leverages background semantics to suppress false positives and utilizes target localization cues to refine segmentation boundaries. This joint formulation facilitates a more holistic and robust feature representation, enhancing the discriminative capability of the network in cluttered infrared scenes.
IV-B1 Dynamic Channel Selection Module
Unlike conventional multi-task learning strategies that typically employ static or full feature sharing, which may lead to the propagation of redundant or even conflicting information across tasks, we introduce a dynamic cross-task feature exchange mechanism. This mechanism adaptively selects and exchanges task-relevant channel-wise features between the detection and segmentation branches. The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution. This dynamic and selective interaction not only strengthens the synergy between detection and segmentation but also effectively mitigates the risk of negative transfer, thereby enhancing overall detection performance in complex infrared environments.
@@ IV-B1 Dynamic Channel Selection Module
Projection Module. Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we design a Projection Module to generate a channel-wise attention vector $\mathbf{m}\in\mathbb{R}^{C}$. This design is inspired by our previous work, Large Selective Kernel Network (LSKNet) [23], and aims to enhance channel discrimination by integrating local feature extraction with global context modeling.
This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion.
  • local feature extraction: To capture fine-grained local representations, we first apply two independent convolutional operations to the input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, each reducing the channel dimension to $C/2$. These two branches are designed to extract complementary local patterns.
    $\mathbf{X}_{1}$ $=\text{Conv}_{1}(\mathbf{X}),$ (5)
    $\mathbf{X}_{2}$ $=\text{Conv}_{2}(\mathbf{X}),$
    $\mathbf{X}_{c}$ $=\mathbf{X}_{1}\oplus\mathbf{X}_{2},$
    where $\oplus$ denotes concatenation operation, resulting in $\mathbf{X}_{c}\in\mathbb{R}^{C\times H\times W}$.
  • channel-wise statistics computation: To integrate global context and emphasize informative regions, we apply both average pooling and max pooling along the channel dimension to form a compact spatial descriptor, followed by a convolutional projection and sigmoid activation.
    $\text{Feat}=\sigma(\mathcal{S}(\text{avg}(\mathbf{X}_{c})\oplus\max(\mathbf{X}_{c}))),$ (6)
@@
    where $\max$ and avg are the maximum and average pooled spatial feature descriptors. $\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features, and $\sigma(\cdot)$ is the sigmoid function. The obtained $\text{Feat}\in\mathbb{R}^{2\times H\times W}$ is a tensor, which represents the weighted fusion of $\max(\mathbf{X})$ and $\text{avg}(\mathbf{X})$.
  • weighted fusion: The attention map Feat is split along the channel axis and used to reweight the previously extracted features $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ through element-wise multiplication. The reweighted features are then concatenated to form the output of the projection module:
    $\mathbf{Z}$ $=\mathcal{F}(\mathbf{X}_{1},\text{Feat})\oplus\mathcal{F}(\mathbf{X}_{2},\text{Feat}),$ (7)
    $\mathbf{m}$ $=\text{mean}\left(\sigma\left(\mathbf{X}\odot\mathbf{Z}\right)\right),$
    where $\odot$ denotes element-wise multiplication, and the operator $\mathcal{F}(\mathbf{X}_{i},\text{Feat})$ denotes a channel-wise spatial reweighting operation that modulates the local feature $\mathbf{X}_{i}$ using the corresponding attention map $\text{Feat}_{i}$. mean denotes a global average pooling operation that aggregates feature representations across spatial dimensions, reducing spatial information into a channel-wise global descriptor. $\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels. Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
@@
where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample. ${I}_{\text{topk}}\in\mathbb{N}^{K}$ represents the corresponding channel indices for these weights. This dynamic selection mechanism enables adaptive identification of the most informative channels for each sample, thereby enhancing the effectiveness of cross-task feature interaction.
@@
TABLE II: Comparison with Other State-of-the-art object detection methods on DenseSIRST.
Method Backbone FLOPs Params DenseSIRST
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
                         End2End
DETR [24] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
Sparse R-CNN [25] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
Deformable DETR [26] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
DAB-DETR [28] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
@@
One-stage
SSD [36] VGG16 87.552G 23.746M 0.211 0.421 0.178 0.424
RetinaNet [37] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
YOLOv3 [38] Darknet 50.002G 61.949M 0.233 0.424 0.207 0.413
CenterNet [39] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
FCOS [22] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
ATSS [40] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
CentripetalNet [41] HourglassNet 0.491T 0.206G 0.244 0.259 0.201 0.244
AutoAssign [42] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
GFL [43] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
PAA [44] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
VFNet [45] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
PVT-T [46] PVT 41.623G 21.325M 0.109 0.481 0.093 0.501
YOLOF [47] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
YOLOX [48] CSPDarknet 8.578G 8.968M 0.210 0.341 0.180 0.331
TOOD [7] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
DyHead [49] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
DDOD [50] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
RTMDet [51] CSPNeXt 51.278G 52.316M 0.229 0.349 0.212 0.350
EfficientDet [52] EfficientNet 34.686G 18.320M 0.146 0.464 0.094 0.517
IV-B2 Cross-Task Feature Hard-Exchange
The introduction of feature hard-exchange is fundamentally driven by the limitations of conventional feature fusion methods in infrared small target detection. Soft fusion techniques such as element-wise addition or concatenation often compromise discriminative power by diluting sparse target features through weighted averaging, while simultaneously permitting gradient conflicts between detection and segmentation objectives during backpropagation. Furthermore, the inherent semantic gap between heterogeneous tasks—object detection versus background segmentation—causes modality mismatch when transferring unadapted features directly. Feature hard-exchange directly addresses these challenges by enabling replacement of the most discriminative channels identified by DCS Module, ensuring explicit semantic transfer without feature dilution, maintaining gradient isolation through binary masking, and resolving cross-modal discrepancies via feature adaptation.
The cross-task feature hard-exchange mechanism operates through three sequential steps.
  • Exchange Mask Generation: To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width. Using channel indices $I_{\text{topk}}$ from the DCS module (Eq. 8), we construct a binary exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
    where $i$ denotes channel index. This mask $\mathbf{e}$ identifies feature channels for exchange. The 1D channel mask is broadcast to spatial dimensions:
    $\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (10)
    where $\otimes$ indicates broadcasting operation, and $\mathbf{1}_{H\times W}$ is an all-one matrix. Detection and segmentation branches generate $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively.
  • Feature Exchange and Adaptation: Bidirectional hard exchange is performed as:
Compared to soft fusion techniques, hard-exchange provides distinct advantages: complete channel replacement preserves feature integrity by avoiding partial feature loss; gradient isolation ensures that gradients propagate exclusively to the source task for exchanged channels; and binary masks enable computationally efficient zero-parameter exchange. This integrated design maintains high discriminative power in complex infrared scenarios while ensuring feature compatibility through the adapter $\mathcal{A}$.
TABLE III: Comparison with Other State-of-the-art object detection methods on IRSTD-1k and IRSDSS.
Method Backbone FLOPs Params IRSTD-1k IRSDSS
mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$